OpenStudy (kevin8767):

Solve for k. 8k + 2m = 3m + k A k = 70m B k = 7m C k = 7/m D k = m/7

OpenStudy (anonymous):

8k + 2m = 3m + k => 8k + 2m - k = 3m => 7k + 2m = 3m => 7k = 3m - 2m => 7k = 1m => k = 1/7*m => k = m/7

OpenStudy (kevin8767):

ty